l cricket LTE 6:45 PM 1% Ch 7 Hmwk- Break-Even Analysis Make s Submission Grade Detail answer Homework Exercise: For the past ten years, you've worked at a PETCO Salon as a dog groomer. You're thinking of starting your own dog grooming business. You found a place you could rent that's right next to a popular shopping center, and two of your friends (who are also dog groomers) have agreed to work for you. The problem is that you need to borrow money to start the business and your banker has asked for a breakeven analysis. You have prepared the following cost estimates for your first year of operations: Fixed Costs $105,000 Salaries Rent and utilities $36,000 Dashboard Calendar To Do Notifications Inbox
l cricket LTE 6:45 PM 1% Ch 7 Hmwk- Break-Even Analysis Rent a Submission Grade Detail Advertising $2,000 $3,000 Equipment Variable Cost per Dog Shampoo $2.00 Coat conditioner $1.50 Pet cologne $0.75 $1.25 Dog treats $0.50 Hair ribbons You went online and researched grooming prices in your area. Based on your review, you have decided to charge $32 for each grooming. Calendar Dashboard To Do Notifications Inbox
l cricket LTE 6:45 PM 1% Ch 7 Hmwk- Break-Even Analysis have d groom Submission Grade Detail 1. Part 1 1. What's the breakeven point in units -how many dogs will you need to groom in the first year to break even? 2. If you and your two employees groomed dogs five days a week, seven hours a day, fifty weeks a year, how many dogs would each of you need to groom each day? Is this realistic given that it takes one hour to groom a dog? 2. Part 2: 1. If you raised your grooming fee to $38, how many dogs would you need to groom to break even? 2. At this new price, how many dogs will each of you have to groom each day (assuming, again, that the three of you groom dogs fifty weeks a Dashboard Calendar To Do Notifications Inbox
l cricket LTE 6:45 PM 1% Ch 7 Hmwk - Break-Even Analysis veal. how maV QIS WOLK eacio Submission Grade Detail Teaistic givemtiatittakes une our to groom a dog? 2. Part 2 1. If you raised your grooming fee to $38, how many dogs would you need to groom to break even? 2. At this new price, how many dogs will each of you have to groom each day (assuming, again, that the three of you groom dogs fifty weeks a year, five days a week, seven hours a day)? 3. Part 3: 1. Would you start this business? Why/why not? 2. What price would you charge to groom a dog? 3. How could you lower the breakeven point and make the business more profitable? Dashboard Calendar To Do Notifications Inbox
Answer #1

1) Fixed Costs: 105,000 + 36,000 + 2,00 + 3,000 (adding the fixed costs salaries, rent, utilities, advertising , equipment)

= $ 146,000 (assuming this will be the overall fixed cost for year 1)

Variable Costs = 2 + 1.5 + 0.75 + 1.25 + 0.5 = $6 (adding all the variable costs )

Now To break even in year one for example we groom X number of dogs for $32 each

=> X (32-6) = 146,000

=> X= 5615.4 we round off to the next figure i.e. 5616 dogs in year 1

2)

Time to groom one dog = 1 hour

Total Workers = 3 (me and two employees)- considering myself as an employee too

7 hours a day

5 days a week

50 weeks a year

Number of Hours working per employee a year = 7 * 5 * 50 = 1750 per employee

Total for three employees = 52,50 dogs

We need to have more 366 dogs to be groomed i.e. 122 more dogs per employee ==> we need to charge higher per dog grooming. by this number, it doesn't look realistic to work beyond 7 hours a day

3) New Price $ 38 per grooming of a dog

if Y is the total dogs per year

=> Y (38-6) = 146,000

Y = 4563 (rounding off from 4562.5 to next figure)

4) Again at the same schedule, we would groom 52,50 dogs total

If total dogs to be groomed is 4563 we would need to groom 1521 dogs each per year

5) Starting this business seems feasible by charging above $40 per grooming a dog. Hence I would start the business because at $38, each employee would have to work 6 hours a day only to go break even in a year.

6) I would charge $ 40 per dog

7) We could lower the break-even point by reducing the fixed costs

I personally think salaries are quite high if we consider $105,000 for two employees only (considering me as the owner takes no salary but only equities)

Each Employee makes $210 per day which is the grooming price for 5 dogs charged $42 per dog

(we can make them work longer or reduce pay)